Solve the problem below using the most suitable formula with the given values. Show your solution.

A 5.73-L flask at 25 °C contains 0.0388 mol of N2 0.147 mol of CO and 0.0803 mol of H2 . What is the total pressure in the flask in atmosphere?
1
Expert's answer
2021-10-19T09:37:03-0400

Solution.

"p = p(N2)+p(CO)+p(H2)"

"p(X) = \frac{n \times R \times T}{V}"

"p = \frac{R\times T}{V}(n(N2) + n(CO) + n(H2))"

"p = \frac{8.31 \times (25+273.14)}{5.73*10^{-3}}(0.0388 + 0.147 + 0.0803) = 115056.6 \ Pa = 1.14 \ atm"

Answer:

p = 1.14 atm